# SimSort: A Powerful Framework for Spike Sorting by Large-Scale Electrophysiology Simulation

### v1

@@ Abstract
Spike sorting is an essential process in neural recording, which identifies and separates electrical signals from individual neurons recorded by electrodes in the brain, enabling researchers to study how specific neurons communicate and process information. Although there exist a number of spike sorting methods which have contributed to significant neuroscientific breakthroughs, many are heuristically designed, making it challenging to verify their correctness due to the difficulty of obtaining ground truth labels from real-world neural recordings. In this work, we explore a data-driven, deep learning-based approach. We begin by creating a large-scale dataset through electrophysiology simulations using biologically realistic computational models. We then present SimSort, a pretraining framework for spike sorting. Remarkably, when trained on our simulated dataset, SimSort demonstrates strong zero-shot generalization to real-world spike sorting tasks, significantly outperforming existing methods. Our findings underscore the potential of data-driven techniques to enhance the reliability and scalability of spike sorting in experimental neuroscience.
Machine Learning, ICML
@@ 1 Introduction
Understanding the complex computations performed by the brain requires insight into the activity of individual neurons (Lewicki, 1998; Buzsáki, 2004), which is crucial for exploring how information is encoded, processed, and transmitted within neural circuits, as well as decoding the brain’s functional dynamics (Quian Quiroga & Panzeri, 2009; Joshi et al., 2023; Sarkar, 2024). Recent advances in neural recording technologies have made it possible to capture the activity of neurons across multiple regions of the brain with high spatial and temporal precision (Chung et al., 2017; Steinmetz et al., 2021; Chung et al., 2019; Hong & Lieber, 2019). Notably, extracting meaningful information from recording data relies on a critical step known as spike sorting.
Spike sorting is the process of extracting and identifying neural activity from extracellular recordings. It involves two main steps (Fig. 1): spike detection, which extracts spike events from background noise, and spike identification, which assigns these detected spikes to individual neurons (Lefebvre et al., 2016). Spike sorting is indispensable for transforming raw electrical signals into interpretable data that reveal the firing patterns of individual neurons. Accurate spike sorting is essential for linking neural activity to behavior, understanding the functional organization of neural circuits (Sibille et al., 2022), and uncovering mechanisms underlying various sensory and cognitive processes (Liu et al., 2022). Furthermore, its role extends to translational applications, such as improving neural prosthetics and developing closed-loop BCIs, where spike sorting is critical for achieving precise neural decoding and control (Park et al., 2017; Han et al., 2024).
For a long time, spike sorting predominantly relied on heuristic statistics and machine learning approaches (Lewicki, 1998; Quiroga et al., 2004; Rossant et al., 2016; Pachitariu et al., 2016; Chung et al., 2017; Hilgen et al., 2017; Pachitariu et al., 2024). While these methods have promoted neuroscience research, they exhibit several key limitations. First, their sorting results are sensitive to parameter settings and post-processing, which depend on the experimenter’s expertise and must be customized for each dataset. Moreover, these approaches lack a data-driven foundation, limiting their scalability and adaptability across diverse experimental settings, particularly in low signal-to-noise ratio (SNR) and high-variability scenarios, where their performance declines significantly. Recently, deep learning-based methods, such as YASS (Lee et al., 2017) and CEED (Vishnubhotla et al., 2023), have attempted to improve spike sorting by adopting data-driven approaches to enhance detection and clustering accuracy. However, their generalizability and practical use are hindered by the limited training datasets.
@@
Therefore, in this work, we first generated a large-scale labeled dataset to address the scarcity of ground-truth data for spike sorting. Building on this, we present SimSort, a framework that utilizes data-driven approaches to enable fully automated spike sorting. By pretraining a spike detection model on the large-scale dataset, SimSort achieved significant improvements in detection accuracy and adaptability compared to commonly used threshold-based methods. Additionally, SimSort incorporated contrastive learning to enhance waveform feature representations, ensuring robustness against noise. We evaluated SimSort through zero-shot experiments on publicly available datasets without fine-tuning, demonstrating its state-of-the-art performance in spike sorting tasks compared to existing methods.
@@
1) A publicly available, large-scale labeled dataset to address the scarcity of ground-truth data, enabling the development and evaluation of advanced spike sorting methods.
@@
3) A practical and robust tool designed to simplify spike sorting for neuroscientists, offering full automation without reliance on manually defined parameters.
Figure 1: Illustration of the spike sorting pipeline, which consists of two main steps: spike detection and spike identification. For spike detection, typical spike sorting algorithms (left) utilize threshold-based detector relies on fixed voltage thresholds on each channel, and use non-learning method like performing PCA-based clustering on concatenated waveforms for spike identification. Those approaches are sensitive to noise and require manual parameter tuning. In contrast, our proposed framework SimSort (right) use a neural network-based detector replaces the thresholding method for spike detection, enhancing robustness and generalization. For spike identification, the feature embeddings of multi-channel waveforms learned from contrastive learning to improve clustering accuracy.
@@ 2 Preliminaries
The goal of spike sorting is to extract single-neuron spiking activity from extracellular recordings, where the activities of an unknown number of neurons are mixed together. Mathematically, spike sorting is a blind source separation problem (Buccino et al., 2022) and can be formally defined as follows: Let $\mathbf{V}\in\mathbb{R}^{T\times C}$ represent the extracellular recording (voltage) across $C$ electrode channels over $T$ time points. The goal is to infer a set of spike times $\{t_{k}\}_{k=1}^{K}$ and corresponding neuron labels $\{y_{k}\}_{k=1}^{K}$, where $y_{k}\in\{1,\dots,N\}$ represents the neuron identity, and $N$ is the total number of detected neurons.
@@ 3 Related Work
Early spike sorting pipelines typically used thresholds to detect spike events, followed by dimensionality reduction (e.g., PCA or $t$-SNE) and clustering (Pearson, 1901; van der Maaten & Hinton, 2008; Veerabhadrappa et al., 2020). Although this pipeline was straightforward, it often suffered from noise susceptibility, inaccurate detection of low SNR events, and reliance on manual parameter tuning.
Deep learning approaches have been adopted to replace or enhance certain steps in traditional pipelines, aiming to improve robustness, automation, and adaptability to diverse experimental conditions. Autoencoders have been explored as a method for dimensionality reduction (Baldi, 2012; Wu et al., 2019; Eom et al., 2021). YASS (Lee et al., 2017) employed a convolutional neural network for spike detection and waveform cleaning, thereby mitigating clustering errors caused by distorted waveforms. CEED (Vishnubhotla et al., 2023) applied a contrastive learning framework to enforce invariances to amplitude fluctuations, noise, and channel subset changes in the extracted embeddings. However, these methods were limited by their reliance on restricted training datasets, potentially reducing their generalization across diverse experimental conditions.
@@
Given these limitations, we propose a data-driven approach to develop a fully automated spike sorting pipeline that does not rely on manually defined parameters, demonstrating improved generalization across diverse datasets and recording conditions.
@@ 4 Methods
4.1 Dataset Creation
To overcome the challenges of insufficient labeled data, we generated a simulated dataset designed to support robust model training and evaluation.
Simulation
Our simulated dataset was generated with multi-compartmental neuron models from the Blue Brain Project (BBP) Neocortical Microcircuit Models Database (Hay et al., 2011), and it contained 206 neuron models from the juvenile rat somatosensory cortex (layers 1-6), covering 30 distinct neuronal types.
These models provided a comprehensive representation of neuronal diversity with precise morphological and ion channel characteristics, enabling realistic biological simulations of electrophysiology responses. Based on these neuron models, we conducted intracellular simulations to compute neuronal transmembrane currents, which were subsequently used to model extracellular potentials. Below, we detail these two stages.
Intracellular Simulation. Intracellular electrophysiological activity was simulated using the NEURON package (Hines & Carnevale, 1997, 2001). Pink noise, scaled to the rheobase (the minimum current needed to elicit an action potential in a neuron), was injected into each neuron’s soma to evoke stochastic action potentials, simulating biologically realistic background noise. The temporal resolution was set to 0.1 ms, and simulations spanned a total duration of 600 seconds. Neurons were initialized with a resting membrane potential of -70 mV and maintained at a physiological temperature of 34°C. Synaptic mechanisms and biophysical properties were dynamically compiled and loaded to ensure compatibility with the NEURON simulation environment.
For each neuron, a multi-compartmental model was used to calculate transmembrane currents (Lindén et al., 2014; Hagen et al., 2015, 2018; Buccino & Einevoll, 2021). The neuron was divided into compartments, each described as an equivalent electrical circuit. The dynamics of the membrane potential in each compartment $n$ were governed by Kirchhoff’s current law (Holmes, 2009), where the sum of all currents entering or leaving a node must equal zero. Considering that extracellular potential changed much slower than ion channel dynamics, we assumed it constant. The temporal evolution of the membrane potential $V_{n}$ in compartment $n$ was given by $g_{n,n+1}(V_{n+1}-V_{n})-g_{n-1,n}(V_{n}-V_{n-1})=C_{n}\frac{dV_{n}}{dt}+\sum_{j}I_{n}^{(j)},$ where $g_{n,n+1}$ and $g_{n-1,n}$ represented the conductances between neighboring compartments, $C_{n}$ was the membrane capacitance, and $\sum_{j}I_{n}^{(j)}$ accounted for ionic currents and any externally applied currents.
Extracellular Simulation. Extracellular potentials were modeled using the volume conductor theory, which links transmembrane currents to extracellular potentials recorded at electrode sites (Brette & Destexhe, 2012; Buzsáki et al., 2012; Einevoll et al., 2013). In each simulation trial, five neurons were randomly selected from the available models and positioned within a 100 × 100 × 100 $\mu$m3 cubic space. These neurons consisted of different types to represent the diversity of cortical microcircuits. A virtual tetrode electrode with four recording sites was placed randomly within the same space to capture extracellular potentials.
@@ 4.1 Dataset Creation
We used signals simulated from BBP layer 1–5 neuron models for training and validation, reserving signals from BBP layer 6 neuron models exclusively for evaluation, ensured that the test set comprises previously unseen neuron types and configurations, allowing us to effectively assess the model’s generalization capabilities.
@@ 4.2 Spike Detection
Figure 3: Diagram of the spike detection model in SimSort.
We utilized a Transformer-based architecture (Vaswani et al., 2017) for spike detection, trained on the simulated continuous signal dataset detailed in Sec. 4.1. To improve robustness and generalization, data augmentation was integrated into the training process, referring to Fig. 3. We employed a binary cross-entropy loss for the spike detection model:
@@
where the binary label $y_{t}$ denoted spike events, $\hat{y}_{t}$ was the predicted probability, $T$ was the length of the trial, and $w_{p}$ was a hyperparameter to reweigh the positive predictions given the general scarcity of spikes.
Data Augmentation
To improve robustness and generalization in spike detection, data augmentation was applied to each input signal segment $V^{\prime}\in\mathbb{R}^{T\times C}$ (Fig. 3). A subset of channels $C^{\prime}\subseteq\{1,2,\dots,C\}$ was randomly selected with a probability $p$. Various augmentations, such as adding noise, amplitude scaling, and temporal jitter, were applied to enhance the diversity of training data. The detailed implementation of these augmentations is provided in Appendix A.
@@ 4.3 Spike Identification
We formulate the spike identification task as embedding each identified spike waveform $X_{i}\in\mathbb{R}^{L\times C}$, the voltage trace (waveform) of spike $i$, into a latent space to learn robust representations using contrastive learning. Then the learned embeddings are clustered into several groups, indicating putative neurons.
The spike identification model was trained on the simulated spike waveform dataset detailed in Sec. 4.1. The model consisted of several key components, as illustrated in Fig. 4 and described as follows.
Figure 4: Diagram of the spike identification model in SimSort.
Contrastive Learning
We leveraged contrastive learning to learn representations that capture the relationships between spike waveforms (Fig. 4). Contrastive learning sought to map neural data into an embedding space where related examples were positioned nearby, while unrelated examples were placed further apart (Dorkenwald et al., 2023; Yu et al., 2024). For each waveform $X_{i}$, we constructed a triplet by selecting $X_{i}^{+}$ with the same label and $X_{i}^{-}$ with a different label. The objective was (Schroff et al., 2015):
$\begin{split}\mathcal{L}_{\text{triplet}}=\max\big{(}0,&\|f(X_{i})-f(X_{i}^{+})\|_{2}^{2}\\ &-\|f(X_{i})-f(X_{i}^{-})\|_{2}^{2}+\alpha\big{)},\end{split}$ (2)
where $f(X_{i})$ denoted the embedding of $X_{i}$, and $\alpha$ was a margin enforcing separation between positive-negative pairs.
Additionally, we uniformly applied data augmentation to each triplet waveforms. The augmentation methods followed those in Sec. 4.2, with adjusted parameters (Appendix A) for better waveform representation, ensuring robustness to noise, amplitude variations, and temporal jitter.
Denoising
The augmented waveform $X_{i}^{\prime}\in\mathbb{R}^{L\times C}$ then underwent a denoising step using Singular Value Decomposition (SVD) (Eckart & Young, 1936; Cunningham & Yu, 2014). The waveform was first reshaped and decomposed as $X_{i}^{\prime}=U\Sigma V^{\top}$, where $U$, $\Sigma$, and $V$ represented the singular vectors and singular values. The first $k$ components were retained to reconstruct the waveform.
Encoder
The denoised waveform $X_{i,\text{recon}}^{\prime}$ was fed into a Gated Recurrent Unit (GRU) encoder (Cho et al., 2014) (Fig. 4), which processed the input sequentially. The final hidden state $h_{T}$ served as the learned representation $f(X_{i})$, encapsulating the spatiotemporal features of the waveform for subsequent clustering tasks.
Clustering
After obtaining the waveform representations, we applied dimensionality reduction using UMAP (McInnes et al., 2020a) before performing clustering (Fig. 4). We used two clustering approaches: a parametric algorithm, Gaussian Mixture Model (GMM) (Dempster et al., 1977), and a non-parametric algorithm, Mean Shift (Cheng, 1995).
@@ 5 Results
With the large-scale dataset and a pretraining framework, it is natural to ask the following research questions (RQ).
@@
RQ3: How important is data size?
In this section, we aimed to address these RQs with comprehensive validation and analysis. We started by presenting a detailed description of the test dataset and the experiment setup in Sec. 5.1. Next, we showed SimSort’s promising spike sorting results on both simulated and real-world data-based benchmarks in Sec. 5.2 (RQ1). Then, we explained the advantages of SimSort with analysis on spike detection and identification, respectively, in Sec. 5.3 (RQ2). Finally, we investigated the impact of data size in Sec. 5.4 (RQ3).

### v2

# SimSort: A Data-Driven Framework for Spike Sorting by Large-Scale Electrophysiology Simulation

@@ Abstract
Spike sorting is an essential process in neural recording, which identifies and separates electrical signals from individual neurons recorded by electrodes in the brain, enabling researchers to study how specific neurons communicate and process information. Although there exist a number of spike sorting methods which have contributed to significant neuroscientific breakthroughs, many are heuristically designed, making it challenging to verify their correctness due to the difficulty of obtaining ground truth labels from real-world neural recordings. In this work, we explore a data-driven, deep learning-based approach. We begin by creating a large-scale dataset through electrophysiology simulations using biologically realistic computational models. We then present SimSort, a pretraining framework for spike sorting. Trained solely on simulated data, SimSort demonstrates zero-shot generalizability to real-world spike sorting tasks, yielding consistent improvements over existing methods across multiple benchmarks. These results highlight the potential of simulation-driven pretraining to enhance the robustness and scalability of spike sorting in experimental neuroscience.
@@ 1 Introduction
Understanding the complex computations performed by the brain requires insight into the activity of individual neurons [1, 2], which is crucial for exploring how information is encoded, processed, and transmitted within neural circuits, as well as decoding the brain’s functional dynamics [3, 4, 5]. Recent advances in neural recording technologies have enabled capturing the activity of neurons across multiple regions of the brain with high spatial and temporal precision [6, 7, 8, 9]. Notably, extracting meaningful information from recordings relies on a critical step known as spike sorting.
Spike sorting is the process of extracting and identifying neural activity from extracellular recordings. It involves two main steps (Fig. 1): spike detection, which extracts spike events from background noise, and spike identification, which assigns these detected spikes to individual neurons [10]. Spike sorting is indispensable for transforming raw electrical signals into interpretable data that reveal the firing patterns of individual neurons. Accurate spike sorting is essential for linking neural activity to behavior, understanding the functional organization of neural circuits [11], and uncovering mechanisms underlying various sensory and cognitive processes [12]. Furthermore, its role extends to translational applications, such as improving neural prosthetics and developing closed-loop BCIs, where spike sorting is critical for achieving precise neural decoding and control [13, 14].
For a long time, spike sorting predominantly relied on heuristic statistics and machine learning approaches [1, 15, 16, 17, 6, 18, 19]. While these methods have promoted neuroscience research, they exhibit several key limitations. First, their sorting results are sensitive to parameter settings and post-processing, which depend on the experimenter’s expertise and must be customized for each dataset. Moreover, these approaches lack a data-driven foundation, limiting their scalability and adaptability across diverse experimental settings, particularly in low signal-to-noise ratio (SNR) and high-variability scenarios, where their performance declines significantly. Recently, deep learning-based methods, such as YASS [20] and CEED [21], have attempted to improve spike sorting by adopting data-driven approaches to enhance detection and clustering accuracy. However, their generalizability and practical use are hindered by the limited training datasets.
@@
Therefore, in this work, we first generated a large-scale labeled dataset to address the scarcity of ground-truth data for spike sorting. Building on this, we present SimSort, a framework that utilizes data-driven approaches to enable fully automated spike sorting. By pretraining a spike detection model on the large-scale dataset, SimSort achieved notable improvements in detection accuracy and adaptability compared to commonly used threshold-based methods. Additionally, SimSort incorporated contrastive learning to enhance waveform feature representations, achieving better robustness against noise. We evaluated SimSort in zero-shot settings on publicly available datasets without fine-tuning, showing its effectiveness in spike sorting tasks compared to existing methods.
@@
1) A publicly available, large-scale labeled dataset to help address the scarcity of ground-truth data, supporting the development and evaluation of learning-based spike sorting methods.
@@
3) An off-the-shelf pretrained model for neuroscientists to use***See https://SimSortTool.github.io for the model, code and usage instructions., enabling fully automated spike sorting on Tetrode recordings without dependence on hand-tuned parameters.
Figure 1: Pipeline of spike sorting, which consists of two main steps: spike detection and spike identification. For spike detection, typical spike sorting algorithms (left) utilize threshold-based detector relies on fixed voltage thresholds on each channel, and use non-learning method like performing PCA-based clustering on concatenated waveforms for spike identification. Those approaches are sensitive to noise and require manual parameter tuning. In contrast, our proposed framework SimSort (right) use a neural network-based detector replaces the threshold method for spike detection, enhancing robustness and generalization. For spike identification, the feature embeddings of multi-channel waveforms learned from contrastive learning to improve clustering accuracy.
@@ 2 Preliminaries
The goal of spike sorting (Figure 1) is to extract single-neuron spiking activity from extracellular recordings, where the activities of an unknown number of neurons are mixed together. Mathematically, spike sorting is a blind source separation problem [22] and can be formally defined as follows: Let $\mathbf{V}\in\mathbb{R}^{T\times C}$ represent the extracellular recording (voltage) across $C$ electrode channels over $T$ time points. The goal is to infer a set of spike times $\{t_{k}\}_{k=1}^{K}$ and corresponding neuron labels $\{y_{k}\}_{k=1}^{K}$, where $y_{k}\in\{1,\dots,N\}$ represents the neuron identity, and $N$ is the total number of detected neurons.
@@ 3 Related Work
Early spike sorting pipelines typically used thresholds to detect spike events, followed by dimensionality reduction (e.g., PCA or $t$-SNE) and clustering [23, 24, 25]. Although this pipeline was straightforward, it often suffered from noise susceptibility, inaccurate detection of low SNR events, and reliance on manual parameter tuning.
Deep learning approaches have been adopted to replace or enhance certain steps in traditional pipelines, aiming to improve robustness, automation, and adaptability to diverse experimental conditions. Autoencoders have been explored as a method for dimensionality reduction [26, 27, 28]. YASS [20] employed a convolutional neural network for spike detection and waveform cleaning, thereby mitigating clustering errors caused by distorted waveforms. CEED [21] applied a contrastive learning framework to enforce invariances to amplitude fluctuations, noise, and channel subset changes in the extracted embeddings. However, these methods were limited by their reliance on restricted training datasets, potentially reducing their generalization across diverse experimental conditions.
@@
To overcome these limitations, we propose a data-driven approach to develop a fully automated spike sorting pipeline that does not rely on manually defined parameters, demonstrating improved generalization across diverse datasets and recording conditions.
@@ 4 Methods
4.1 Dataset Creation
To address the scarcity of labeled spike data, we generated a large-scale synthetic dataset using biophysically detailed neuron models from the Blue Brain Project (BBP) [29], covering 206 neuron models from layers 1–6 of the juvenile rat somatosensory cortex. These models capture diverse morphologies and electrophysiological dynamics across 30 neuronal types.
Electrophysiology Simulation
Based on these models, we first simulated intracellular activity by injecting noise currents into multi-compartment neurons using the NEURON simulator [30], producing realistic spiking responses. Then, extracellular potentials were computed via volume conductor theory, summing transmembrane currents across compartments to obtain virtual recordings from tetrode electrodes randomly placed near the neurons (Figure 2; full technical details in Appendix A).
Each simulation trial included five neurons of varying types placed in a $100\times 100\times 100~{}\mu$m3 volume, with randomized positions for both cells and electrodes. We generated 8,192 trials in total, yielding over 40,000 annotated neuron instances. Ground-truth spike times were extracted from intracellular traces and aligned to extracellular recordings.
@@ 4.1 Dataset Creation
We used signals simulated from BBP layers 1–5 neuron models for training and validation, reserving signals from BBP layer 6 neuron models exclusively for evaluation, ensured that the test set comprises previously unseen neuron types and configurations, allowing us to effectively assess the model’s generalization capabilities.
@@ 4.2 Spike Detection
Figure 3: Spike detection model in SimSort.
We utilized a Transformer-based architecture [31] for spike detection, trained on the simulated continuous signal dataset detailed in Sec. 4.1. For robustness and generalization, data augmentation was integrated into the training process, referring to Fig. 3. We employed a binary cross-entropy loss for the spike detection model:
@@
where the binary label $y_{t}$ denoted spike events, $\hat{y}_{t}$ was the predicted probability, $T$ was trial length, and $w_{p}$ was a hyperparameter to reweigh the positive predictions given the general scarcity of spikes.
Data Augmentation: To improve robustness and generalization in spike detection, data augmentation was applied to each input signal segment $V^{\prime}\in\mathbb{R}^{T\times C}$ (Fig. 3). A subset of channels $C^{\prime}\subseteq\{1,2,\dots,C\}$ was randomly selected with a probability $p$ to apply augmentation. Various augmentations, such as adding noise, amplitude scaling, and temporal jitter, were applied to enhance the diversity of training data. The detailed implementation of these augmentations is provided in Appendix A.
@@ 4.3 Spike Identification
We formulate spike identification as embedding each identified spike waveform $X_{i}\in\mathbb{R}^{L\times C}$, the voltage trace (waveform) of spike $i$, into a latent space to learn robust representations using contrastive learning. Then the learned embeddings are clustered into several groups, indicating putative neurons.
The spike identification model was trained on the simulated spike waveform dataset (Sec. 4.1). The model consisted of several key components, as illustrated in Fig. 4 and described as follows.
Figure 4: Spike identification model in SimSort.
Contrastive Learning: We leveraged contrastive learning to learn representations that capture the relationships between spike waveforms (Fig. 4). Contrastive learning sought to map neural data into an embedding space where related examples were positioned nearby, while unrelated examples were placed further apart [32, 33]. For each waveform $X_{i}$, we constructed a triplet by selecting $X_{i}^{+}$ with the same label and $X_{i}^{-}$ with a different label. The objective was [34]:
$\mathcal{L}_{\text{triplet}}=\max\big{(}0,\|f(X_{i})-f(X_{i}^{+})\|_{2}^{2}-\|f(X_{i})-f(X_{i}^{-})\|_{2}^{2}+\alpha\big{)},$ (2)
where $f(X_{i})$ denoted the embedding of $X_{i}$, and $\alpha$ was a margin enforcing separation between positive-negative pairs. Additionally, we uniformly applied data augmentation to each triplet waveforms. The augmentation methods followed those in Sec. 3, with adjusted parameters (Appendix A) for better waveform representation, ensuring robustness to noise, amplitude variations, and temporal jitter.
Denoising: The augmented waveform $X_{i}^{\prime}\in\mathbb{R}^{L\times C}$ then underwent a denoising step using Singular Value Decomposition (SVD) [35, 36]. The waveform was first reshaped and decomposed as $X_{i}^{\prime}=U\Sigma V^{\top}$, where $U$, $\Sigma$, and $V$ represented the singular vectors and singular values. The first $k$ components were retained to reconstruct the waveform.
Encoder: The denoised waveform $X_{i,\text{recon}}^{\prime}$ was fed into a Gated Recurrent Unit (GRU) encoder [37] (Fig. 4), which processed the input sequentially. The final hidden state $h_{T}$ served as the learned representation $f(X_{i})$, encapsulating the spatiotemporal features of the waveform for subsequent clustering tasks.
Clustering: After obtaining the waveform representation, we applied dimensionality reduction using UMAP [38] before clustering (Fig. 4). We tested two clustering approaches: a parametric algorithm, Gaussian Mixture Model (GMM) [39], and a non-parametric one, Mean Shift (MS) [40].
@@ 5 Results
With the large-scale dataset and a pretraining framework, we try to answer the following research questions (RQ).
@@
RQ3: Does data scale hold the key of effective spike sorting?